OpenStudy (anonymous):

A square pyramid has a base with perimeter of 28 inches and a height of 24 inches. Find the surface area and volume of the pyramid

OpenStudy (saifoo.khan):

Dont u know the formulas?

OpenStudy (anonymous):

no they are not in front of me this is my last question i have 3 minutes left on exam

OpenStudy (anonymous):

can u give me the formulas?

OpenStudy (saifoo.khan):

volume = 1/3 base area * height

OpenStudy (saifoo.khan):

28/ 4 = 7in / side. base area = 7 * 7 = 49

OpenStudy (saifoo.khan):

1/3 * 49 * 24

OpenStudy (saifoo.khan):

volume = 392in^3

OpenStudy (anonymous):

ok thats what i got im doing surface now

OpenStudy (saifoo.khan):

surface area = 2 bs + b^2

OpenStudy (anonymous):

what is bs?

OpenStudy (anonymous):

nvm lol base sum

OpenStudy (saifoo.khan):

lol, base * perimeter i guess

OpenStudy (anonymous):

wait what

OpenStudy (saifoo.khan):

let me see.

OpenStudy (saifoo.khan):

this one is hard. :O

OpenStudy (anonymous):

lol great

OpenStudy (anonymous):

oh well i only have a minute left ill skip it

OpenStudy (saifoo.khan):

dang! this is hard/ :O

OpenStudy (saifoo.khan):

SA = 336

OpenStudy (saifoo.khan):

Got it. :D